# https://arc2face.github.io/

### v1

@@ 2.2 Subject-specific Facial Generation
Closest to ours lies a series of recent works that condition Diffusion Models directly on facial features for tuning-free personalization. Various approaches, including FastComposer [86], PhotoVerse [10] and PhotoMaker [45], employ features from the CLIP [59] image encoder to represent the input subject. However, this representation is constrained by the CLIP’s facial encoding abilities. Methods such as Face0 [80], DreamIdentity [11], and PortraitBooth [56] condition the model additionally on FR embeddings for improved fidelity. Similarly, IP-Adapter [90] uses a decoupled cross-attention mechanism to separate text from subject conditioning, and has released post-publication [91], an impressive model based on ID features. InstantID [83] extends [91] with an additional network for stronger ID guidance and facial landmarks conditioning. Finally, FaceStudio [87] learns combined CLIP and ID-embeddings, achieving impressive stylizations.
@@ 4.1.2 Prompt Selection
Refer to caption (a) Synth-500 Refer to caption (b) AgeDB
@@ 4.1.3 Datasets and Metrics
As depicted in Fig. 5, existing methods, particularly [45, 86] that rely on CLIP features, struggle to preserve the ID without detailed text descriptions of the subject. InstantID [83] attains the highest similarity, behind Arc2Face. However, this results in limited pose/exp. diversity as it further uses the input facial landmarks to constrain generation. Arc2Face does not require any text or spatial conditions and achieves the highest facial similarity as well as diversity and realism across both datasets. This particularly highlights the efficiency of ID-embeddings in the context of face generation against CLIP image or text features. A visual comparison with the above methods is provided in Fig. 6, whereas additional qualitative results and visualizations are provided in the Supp. Material. We further conducted a user study to compare with the second best-performing method in terms of ID similarity, i.e. InstantID [83]. In particular, we asked 50 users to choose the method whose result best resembles the input face, regardless of quality or realism, for a randomly selected set of 30 IDs from our datasets. The two methods were randomly presented side-by-side. Fig 5 confirms a strong preference for Arc2Face in terms of ID resemblance.
@@ 4.4 Ablation Studies
Principal Component Analysis To explore the intrinsic dimensionality of our facial ID representation, we conducted PCA on the ArcFace embeddings derived from WebFace42M images. Fig. 8(a) depicts the cumulative percentage of variance explained by the principal components. Notably, maintaining facial fidelity requires at least 300-400 components, as fewer components result in noticeable distortion, as shown in Fig. 8(b). This reveals the challenge of compressing ArcFace embeddings significantly, emphasizing the inherent complexity of our facial ID representation.
@@
Refer to caption (a) Synth-500 Refer to caption (b) AgeDB
Refer to caption (a) Cumulative Var. Refer to caption (b) PCA projections
@@ 6 Conclusion
Acknowledgements. S. Zafeiriou and part of the research was funded by the EPSRC Fellowship DEFORM (EP/S010203/1) and EPSRC Project GNOMON (EP/X011364/1).

### v2

@@ 2.2 Subject-Specific Facial Generation
Closest to ours lies a series of recent works that condition Diffusion Models directly on facial features for tuning-free personalization. Various approaches, including FastComposer [86], PhotoVerse [10] and PhotoMaker [45], employ features from the CLIP [59] image encoder to represent the input subject. However, this representation is constrained by the CLIP’s facial encoding abilities. Methods such as Face0 [80], DreamIdentity [11], and PortraitBooth [56] condition the model additionally on FR embeddings for improved fidelity. Similarly, IP-Adapter [90] uses a decoupled cross-attention mechanism to separate text from subject conditioning and has released, post-publication, [91] an impressive model based on ID features. InstantID [83] extends [91] with an additional network for stronger ID guidance and facial landmarks conditioning. Finally, FaceStudio [87] learns combined CLIP and ID-embeddings, achieving impressive stylizations.
@@ 4.1.2 Prompt Selection
(a) Synth-500
(b) AgeDB
@@ 4.1.3 Datasets and Metrics
As depicted in Fig. 5, existing methods, particularly [45, 86] that rely on CLIP features, struggle to preserve the ID without detailed text descriptions of the subject. InstantID [83] attains the highest similarity, behind Arc2Face. However, this results in limited pose/exp. diversity as it further uses the input facial landmarks to constrain generation. Arc2Face does not require any text or spatial conditions and achieves the highest facial similarity as well as diversity and realism across both datasets. This particularly highlights the efficiency of ID-embeddings in the context of face generation against CLIP image or text features. A visual comparison with the above methods is provided in Fig. 6, whereas additional qualitative results and visualizations are provided in the Supp. Material. We further conducted a user study to compare with the second best-performing method in terms of ID similarity, i.e. InstantID [83]. In particular, we asked 50 users to choose the method whose result best resembles the input face, regardless of quality or realism, for a randomly selected set of 30 IDs from our datasets. The two methods were randomly presented side-by-side. Fig. 5 confirms a strong preference for Arc2Face in terms of ID resemblance.
@@ 4.4 Ablation Studies
Principal Component Analysis To explore the intrinsic dimensionality of our facial ID representation, we conducted PCA on the ArcFace embeddings derived from WebFace42M images. Fig. 8(c) depicts the cumulative percentage of variance explained by the principal components. Notably, maintaining facial fidelity requires at least 300-400 components, as fewer components result in noticeable distortion, as shown in Fig. 8(d). This reveals the challenge of compressing ArcFace embeddings significantly, emphasizing the inherent complexity of our facial ID representation.
(a) Synth-500
(b) AgeDB
@@
(c) Cumulative Var.
(d) PCA projections
@@ 6 Conclusion
Acknowledgements. S. Zafeiriou and part of the research was funded by the EPSRC Fellowship DEFORM (EP/S010203/1) and EPSRC Project GNOMON (EP/X011364/1). B. Kainz received support from the ERC, project MIA-NORMAL 101083647, DFG 512819079, and by the State of Bavaria (HTA). HPC resources were provided by the Erlangen National High Performance Computing Center (NHR@FAU) of the Friedrich-Alexander-Universität Erlangen-Nürnberg (FAU) under the NHR project b180dc. NHR@FAU hardware is partially funded by the German Research Foundation (DFG) – 440719683.